OpenStudy (marcushooks):

Brenton’s weekly pay, P(h) , in dollars, is a function of the number of hours he works, h. He gets paid $20 per hour for the first 40 hours he works in a week. For any hours above that, he is paid overtime at $30 per hour. He is not permitted to work more than 60 hours in a week.

OpenStudy (dumbcow):

\[P(h) =\left(\begin{matrix}20h , 0<h \le 40 \\ 20h+10(h-40), 40 <h \le 60\end{matrix}\right)\]

OpenStudy (marcushooks):

thanks

OpenStudy (marcushooks):

@dumbcow